A bag is filled with yellow and green marbles. There are 74 marbles in the bag. If there are 10 less yellow marbles than green marbles, how many yellow marbles are in the bag?
Answer #1

Type 1 Given Total two type of marbles = 74 Yellow marble Type 2 Green marble Let T = Number of yellow marble Y = Number of GAgain put the value of y in equation 2 x=y-10 = x=42-10 x=32 So Yellow marbles 32 Green marbles = 42